1. The length of the shadow of a flagpole was found to be 72 feet. The shadow of a 3 foot picket fence in line with the flagpole was 4 feet. What is the height of the flagpole?

Answer by checkley79(3341)

3/4=X/72 CROSS MULTIPLY.
4X=3*72
4X=216
X=216/4
X=54 FEET HIGH FOR THE FLAG POLE.